A system consists of a sealed copper tank whose mass is 20kg, which contains 5 kg of liquid water, and an electrical register (negligible mass). The system is insulated on its outer surface. Initially, the temperature of the copper is 27˚C and the temperature of the water is 50˚C. The electric resister transfers 100 kJ of energy to the system. Eventually the system comes to equilibrium.

a) Sketch the system. Clearly label all components and include arrows indicating work or heat transfer processes.

b) Determine the final equilibrium temperature, in ˚C.
Answer #1

from Data Book Given Date : Mars of Cu Tanie = 20 kg Mais of water in it = 5kg ( specific heat) Gee = 0.385 ks/19k (specificAccording to first law of Thermodynamic for a process: ISO = du + St / S8 = (Ug-()+dh where, so = Heat Interaction et systemSQzo ( since Insulated from outside) Ug = (y cu + (la) water = (mCT u + (mcTf) weser ={(Mc) out (MC) water} TJ (1. u, = (u) c=) Put abore Value in gut law Eq for Process : 0 = {(mc leu + (Chater} } - {{ mC Til au foc i weser} 100 0 = {(2010:385) +(5